1. A perfectly competitive firm sells its product for $360/unit and has an average total cost function given by: ATC(Q) = 1000/Q + 30 + 1.5Q.

a. What are this firm’s fixed costs? Explain.

b. Determine this firm’s profit maximizing level of output.

c. Calculate this firm’s profits.

2. A perfectly competitive firm sells its product for $200/unit and has a total cost of production given by: C(Q) = 1500 + 40Q+5Q2 .

a. What are this firm’s fixed costs? Explain.

b. Determine this firm’s profit maximizing level of output.

c. Determine this firm’s profits.

d. Should this firm shut down immediately? Explain.
Answer #1

p=360 AC= 1000, 30+ 1. 50 0 6) TC=1000+300+1.502 freins fixed cost=1000 Prefit is mak when P= MR=MC MC- 30+30. P=360 thus 30+